Question 1 If the number of units sold increases (quantity increases), which of the following will NOT change? Sales O VariabQuestion 2 Kellys Karts, LLC reported the following data: Sales quantity 1,000 units; Sales $750 per unit; Variable costs $4Question 3 2 pt Kellys Karts, LLC reported the following Contribution Margin Income Statement. What is the effect on Net IncIf monthly advertising is increased by $10,000, which of the following are affected? Fixed costs Variable costs O Sales ContrKellys Karts, LLC reported the following Contribution Margin Income Statement. What is the effect on Net Income if the sales
Answer #1

QUESTION 1

If the number of units sold increases (quantity increases), FIXED COSTS will not change.

  • Sales will increase as the units sold increase. (provided selling price is not decreased)
  • Variable costs are those costs that vary with the level of output. For every unit sold, variable cost is incurred.
  • Contribution margin = Sales - Variable cost. Therefore, it will also change if the number of units sold is increased.

QUESTION 2

Beak-even in sales dollars: $562,500

  • Breakeven point (value in $) = Fixed costs ÷ Contribution margin ratio
    • Fixed costs = $258,750
    • Contribuion margin ratio = Contribution margin ÷ Sales
      • Sales = $750 × 1,000 units = $750,000
      • Variable costs = $405 × 1,000 units = $405,000
      • Contribution margin = Sales - Variable costs = $750,000 - $405,000 = $345,000
      • Contribuion margin ratio = $345,000 ÷ $750,000 = 0.46
    • Break-even sales in dollars = $258,750 ÷ 0.46 = $562,500

Check:

The breakeven point (BEP) is the volume at which there is neither profits nor losses.

Break-even sales quantity (in units) = Fixed costs ÷ Contribution margin per unit = $258,750 ÷ ($750-$405) = 750 units

Sales (break-even) Less: variable costs Contribution margin Less: fixed costs Profit 750 units Ⓡ$750 $ 562,500 750 units $405

QUESTION 3

The effect on Net Income if the sales quantity increases by 100 units: $3,200 increase

Since all the fixed costs are already recovered and the firm is earning profits, the increase in profit = increase in quantity × contribution per unit = 100 units × $32 = $3,200

Check:

Sales (break-even) Less: variable costs Contribution margin Less: fixed costs net income 800 units * $80 800 units x $48 750

  • Fixed costs do not change with change in units sold upto a certain extent. Therefore, it is constant at $12,000

If the monthly advertising is increased by $10,000, fixed costs are affected.

The effect on Net Income if

  • sales quantity increases 200 units
  • sales price decreases $10
  • monthly advertising expense increases by $5,000 (fixed cost)
    • $12,000 decrease

Sales (break-even) Less: variable costs Contribution margin Less: fixed costs Net loss (1000 units + 200 units) x ($60 - $10)